StatChapter03 Part 2: Problem 6 30 13 43 63 37 100 Answer
Answer #1

Let F= female ,      M = Male , D = college degree       N = No college degree

From given table,

a) Consider,

         330.33 P(Fn D)100

         63 P(F) 0.63

         PD-0.57 00 0.57

Now, By Addition theorem

         P(Fcup D)=P(F)+P(D)-P(Fcap D)

                              0.63 + 0.57-0.33

                              =0.87

   Hence, Answer : 0.87

b) Similarly, From table

          370.37 P(M)100

           P(N)=rac{43}{100}=0.43

          13 100

   By Addition theorem,

           P(Mcup N)=P(M)+P(N)-P(Mcap N)

                                 =0.37+0.43-0.13

                                =0.67

   Hence, Answer : 0.67

c) From Table

              30 100

              P(F) = 0.63

              P(N)=0.43

    By Addition theorem,

               P(Fcup N)=P(F)+P(N)-P(Fcap N)

                                    =0.63+0.43-0.30

                                    =0.93

    Hence, Answer : 0.93